Dlikhtar757:

C= 5 9 (F−32) The equation above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5 9 degree Celsius. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. A temperature increase of 5 9 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A) I only B) II only C) III only D) I and II only

Vocaloid:

Fixing some formatting: C= (5/9)(F−32) The equation above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? statement I: A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of (5/9) degree Celsius. statement II: A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. statement III: A temperature increase of (5/9) degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius.

Vocaloid:

evaluate each statement one by one. statement I: A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of (5/9) degree Celsius. ---> take a look at the equation C= (5/9)(F−32). when you increase F by 1, how much does C increase by? you can try treating this equation as a y = mx + b form, where m is the rate of change, and thus the amount that C increases when F increases by 1. statement II: to get Fahrenheit in terms of Celcius, solve the equation for F. from there, see how F changes when you increase C by 1.8. to do this, you can set C = 0, calculate F, and then change C to 1.8 and see how much F changes. statement III: looking at your equation from statement II, how much does F change when C changes by 1? use the same logic as you did from the earlier parts. once you've determined which statement(s) is/are true, see which answer choice corresponds.